Betty is now three times as old as mike,but in five years she will be only twice as old as mike. how old is mike right now

Found 2 solutions by stanbon, Natolino1983:
Answer by stanbon(75887)

Betty is now three times as old as mike,but in five years she will be only twice as old as mike. how old is mike right now
--------
b = 3m
b+5 = 2(m+5)
--------------------
Substitute for "b" and solve for "m"::
3m+5 = 2m+10
m = 5 yrs (mike's age now)
-------------
Cheers,
Stan H.
-------------

Answer by Natolino1983(23)

Let B: Betty's age (today)
M: Mike's age (today).
We have the equations:
1) B= 3×M
2)(B +5) = 2×(M+5)
So 1) in 2) 3×M + 5 = 2×M + 10
M = 5.
ANSWER: Mike is 5 years-old ( Betty is 15 years-old).
Follow me on Twitter for more @natolino_